# Renormalization of the three-flavor quark-meson diquark model

### v1

@@ I Introduction
In Refs. [20, 21], we proposed the two-flavor quark-meson diquark (QMD) model as a renormalizable low-energy effective model to describe the 2SC phase of dense QCD (See also Ref. [22]). The starting point is the quark-meson model (QM) which is used to describe normal quark matter. Describing the color superconducting phases requires the introduction of diquark effective degrees of freedom in the quark-meson model. The diquarks interact with mesons and quarks, and with themselves. The addition of the diquarks gives rise to a few new terms in the low-energy effective Lagrangian. The allowed terms are dictated by symmetry and renormalizability. The couplings of these terms are a priori unknown, and we consider them free. In Refs. [20, 21], it was found that the results are sensitive mainly to the diquark mass parameter $m_{\Delta}$ and the quark-diquark coupling $g_{\Delta}$. By choosing $m_{\Delta}=500-900$ MeV and $g_{\Delta}\approx g$, where $g$ is the Yukawa coupling, we obtained a transition to 2SC matter at around $\mu_{B}=300$ MeV and a gap of 50-150 MeV in the range of baryon chemical potentials where the 2SC is believed to exist. Finally, it was shown that in the limit $\mu\rightarrow\infty$, the gap approaches a constant and the speed of sound relaxes to the conformal limit from above. In Refs. [20, 21], the sigma and the three pions were included as mesonic degrees of freedom. In this case, the model is equivalent to the $O(4)$ linear sigma model. This model does not allow for different light quark masses, and therefore no explicit isospin symmetry breaking. Even if the light quarks are degenerate in the vacuum, isospin can be broken at finite density due to non-zero chemical potentials. For example, in the 2SC phase with charge neutrality imposed, the isospin chemical potential is non-vanishing, and might lead to $m_{u}\neq m_{d}$. Therefore, we first consider the two-flavor system and show how this feature can be incorporated into the model by extending the meson sector.
@@ II Two-flavor quark-meson diquark model
$\mu_{ur}$ $=$ $\mu_{ug}=\mu-{2\over 3}\mu_{e}+{1\over 2}\mu_{3}+{1\over 3}\mu_{8}\;,$ (12)
$\mu_{dr}$ $=$ $\mu_{dg}=\mu+{1\over 3}\mu_{e}-{1\over 2}\mu_{3}+{1\over 3}\mu_{8}\;,$ (13)
$\mu_{sr}$ $=$ $\mu_{sg}=\mu+{1\over 3}\mu_{e}+{1\over 3}\mu_{8}\;,$ (14)
$\mu_{ub}$ $=$ $\mu-{2\over 3}\mu_{e}+{1\over 2}\mu_{3}-{2\over 3}\mu_{8}\;,$ (15)
$\mu_{db}$ $=$ $\mu+{1\over 3}\mu_{e}-{1\over 2}\mu_{3}-{2\over 3}\mu_{8}\;,$ (16)
@@
In this form, the invariance under $SU(3)_{2}\times SU(3)_{2}$ transformations is evident. Moreover, the coefficient of the symmetry breaking term is generalized to
@@
$\delta{\cal L}_{\rm scalar}$ $=$ $c\left[\det\Sigma+\det\Sigma^{\dagger}\right]-{\lambda_{2}}\langle(\Sigma^{\dagger}\Sigma)^{2}\rangle$ (30)
$-\lambda_{\rm det}\left[\det\Sigma+\det\Sigma^{\dagger}\right]\langle\Sigma^{\dagger}\Sigma\rangle\;.$
The determinant term is induced by instantons and is invariant under $SU(2)_{L}\times SU(2)_{R}$. It equals $\sigma^{2}+\vec{\pi}^{2}-\eta^{2}-\vec{a}^{2}$. Without the new fields, this term is equal to $\langle\Sigma^{\dagger}\Sigma\rangle$ and is therefore not an independent operator. There is also another independent quartic term, namely $\langle(\Sigma^{\dagger}\Sigma)^{2}\rangle=(\sigma^{2}+\vec{\pi}^{2})^{2}+(\eta^{2}+\vec{a}^{2})^{2}+(\sigma^{2}+\vec{\pi}^{2})(\eta^{2}+\vec{a}^{2})$. The last term Eq. (30) is necessary if we consider the pion condensed phase at finite $\mu_{I}$ and the resulting isospin breaking. There is also a new interaction term between the scalars/pseudoscalars and the diquarks, namely the product of the building blocks from each sector,
@@ V Renormalization of $\Omega$ in the three-flavor CFL phase
which shows that $c_{s}^{2}$ approaches the conformal limit from above. This is a generic feature of the model: The same behavior was found in the case of finite isospin chemical potential $\mu_{I}$ and pion condensation [37, 38, 39, 21].

### v2

@@ I Introduction
In Refs. [20, 21], we studied in detail the properties of the two-flavor quark-meson diquark (QMD) model as a renormalizable low-energy effective model to describe the 2SC phase of dense QCD, See also Ref. [22, 23, 24]. The starting point is the quark-meson model (QM) which is used to describe normal quark matter. Describing the color superconducting phases requires the introduction of diquark effective degrees of freedom in the quark-meson model. The diquarks interact with mesons and quarks, and with themselves. The addition of the diquarks gives rise to a few new terms in the low-energy effective Lagrangian. The allowed terms are dictated by symmetry and renormalizability. The couplings of these terms are a priori unknown, and we consider them free. In Refs. [20, 21], it was found that the results are sensitive mainly to the diquark mass parameter $m_{\Delta}$ and the quark-diquark coupling $g_{\Delta}$. By choosing $m_{\Delta}=500-900$ MeV and $g_{\Delta}\approx g$, where $g$ is the Yukawa coupling, we obtained a transition to 2SC matter at around $\mu_{B}=300$ MeV and a gap of 50-150 MeV in the range of baryon chemical potentials where the 2SC is believed to exist. Finally, it was shown that in the limit $\mu\rightarrow\infty$, the gap approaches a constant and the speed of sound relaxes to the conformal limit from above. In Refs. [20, 21], the sigma and the three pions were included as mesonic degrees of freedom. In this case, the model is equivalent to the $O(4)$ linear sigma model. This model does not allow for different light quark masses, and therefore no explicit isospin symmetry breaking. Even if the light quarks are degenerate in the vacuum, isospin can be broken at finite density due to non-zero chemical potentials. For example, in the 2SC phase with charge neutrality imposed, the isospin chemical potential is non-vanishing, and might lead to $m_{u}\neq m_{d}$. Therefore, we first consider the two-flavor system and show how this feature can be incorporated into the model by extending the meson sector.
@@ II Two-flavor quark-meson diquark model
$\mu_{ur}$ $=$ $\mu-{2\over 3}\mu_{e}+{1\over 2}\mu_{3}+{1\over 3}\mu_{8}\;,$ (12)
$\mu_{dr}$ $=$ $\mu+{1\over 3}\mu_{e}+{1\over 2}\mu_{3}+{1\over 3}\mu_{8}\;,$ (13)
$\mu_{sr}$ $=$ $\mu+{1\over 3}\mu_{e}+\frac{1}{2}\mu_{3}+{1\over 3}\mu_{8}\;,$ (14)
$\mu_{ug}$ $=$ $\mu-{2\over 3}\mu_{e}-{1\over 2}\mu_{3}+{1\over 3}\mu_{8}\;,$ (15)
$\mu_{dg}$ $=$ $\mu+{1\over 3}\mu_{e}-{1\over 2}\mu_{3}+{1\over 3}\mu_{8}\;,$ (16)
$\mu_{sg}$ $=$ $\mu+{1\over 3}\mu_{e}-\frac{1}{2}\mu_{3}+{1\over 3}\mu_{8}\;,$ (17)
$\mu_{ub}$ $=$ $\mu-{2\over 3}\mu_{e}-{2\over 3}\mu_{8}\;,$ (18)
$\mu_{db}$ $=$ $\mu+{1\over 3}\mu_{e}-{2\over 3}\mu_{8}\;,$ (19)
@@
In this form, the invariance under $SU(2)_{L}\times SU(2)_{R}$ transformations is evident. Moreover, the coefficient of the symmetry breaking term is generalized to
@@
$\delta{\cal L}_{\rm scalar}$ $=$ $c\left[\det\Sigma+\det\Sigma^{\dagger}\right]-{\lambda_{2}}\langle(\Sigma^{\dagger}\Sigma)^{2}\rangle$
$-\lambda_{\rm det}\left[\det\Sigma+\det\Sigma^{\dagger}\right]\langle\Sigma^{\dagger}\Sigma\rangle\;.$ (33)
The determinant term is induced by instantons and is invariant under $SU(2)_{L}\times SU(2)_{R}$. It equals $\sigma^{2}+\vec{\pi}^{2}-\eta^{2}-\vec{a}^{2}$. Without the new fields, this term is equal to $\langle\Sigma^{\dagger}\Sigma\rangle$ and is therefore not an independent operator. There is also another independent quartic term, namely $\langle(\Sigma^{\dagger}\Sigma)^{2}\rangle=(\sigma^{2}+\vec{\pi}^{2})^{2}+(\eta^{2}+\vec{a}^{2})^{2}+(\sigma^{2}+\vec{\pi}^{2})(\eta^{2}+\vec{a}^{2})$. There is also a new interaction term between the scalars/pseudoscalars and the diquarks, namely the product of the building blocks from each sector,
@@ V Renormalization of $\Omega$ in the three-flavor CFL phase
which shows that $c_{s}^{2}$ approaches the conformal limit from above. This is a generic feature of the model: The same behavior was found in the case of finite isospin chemical potential $\mu_{I}$ and pion condensation [39, 40, 41, 42, 21].